Solve. 4(x + 9) = –4
Answer by rmromero(383)


Solve. 4(x + 9) = –4
Use distributive property to simplify the equation
4(x) + 4(9) = -4
4x + 36 = -4
4x + 36 - 36 = -36 - 4 Subtract both sides by 36 to solve for x
4x = -40 Divide both sides by -4
x = -10
Therefore x = -10